OpenStudy (anonymous):

The arc of a circle of radius 1.4cm is equal to ?

OpenStudy (anonymous):

The equation to find the arc length is: \[s = r \theta\]Where s= the arc length, r= the radius, and theta= the angle inside the unit circle. Please note that the angle must be in radians. For example, 30 degree angle would be \[\pi \div6\]and a 60 degree angle would be\[\pi \div3\]To convert angles to radians, just take the angle, mulitply by pi, then divide by 180. Finally simplify.

OpenStudy (anonymous):

UNIT CIRCLE